Consider the R −R2 function r r defined by   (t) = (a) Write down the domain of r (b) Is r (c) Is r continuous at t = 0? continuous at t = 2? (d) Sketch the curve r 26 . .  (t, t2) if t ∈ [−2,0] (t, t) t, t2 if if t ∈ (0,2) t ∈ [2,3]

1
Expert's answer
2022-05-24T14:05:24-0400
"r(t)= \begin{cases}
 (t, t^2) &\text{if } t\in [-2,0] \\
 (t, t) &\text{if } t\in (0,2) \\
 (t, t^2) &\text{if } t\in [2,3] \\
\end{cases}"


  

(a) Domain: "[-2, 3]"


(b)


"r(0)=(0, 0)"

"\lim\limits_{t\to 0^-}r(t)=(0,0)"

"\lim\limits_{t\to 0^+}r(t)=(0,0)"

Then


"\lim\limits_{t\to 0}r(t)=(0,0)=r(0)"

The function "r(t)" is continuous at "t=0."


(c)


"r(2)=(2, 4)"

"\lim\limits_{t\to 2^-}r(t)=(2,2)"

"\lim\limits_{t\to 2^+}r(t)=(2,4)"

"\lim\limits_{t\to 2^-}r(t)\not=\lim\limits_{t\to 2^+}r(t)"

Then


"\lim\limits_{t\to 2}r(t)=\text{does not exist}"

The function "r(t)" is not continuous at "t=2."


(d)